# PAIR-Former: Budgeted Relational Multi-Instance Learning for Functional miRNA Target Prediction

### v1

@@ B.1 Preliminaries: masking view of budgeted execution
Assume B.9, B.10, B.11, B.12 and B.17. Fix any selector $S$ (deterministic or conditioned upon), and consider $\mathcal{H}_{K,S}$. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the draw of $\mathcal{D}$, for all $f\in\mathcal{F}_{K}$,
@@
If additionally B.13 holds with $|\mathcal{S}_{K}|=N_{S}<\infty$, then with probability at least $1-\delta$, uniformly for all $(f,S)\in\mathcal{F}_{K}\times\mathcal{S}_{K}$,
@@
Assume B.11. Let $\mathcal{G}$ be a class of real-valued functions on bags, and define $\ell\circ\mathcal{G}:=\{(B,Y)\mapsto\ell(Y,g(B)):\ g\in\mathcal{G}\}$. Then
@@
Assumption B.17 (Capacity control of $\mathcal{F}_{K}$).
@@
Assume there exists a constant $C_{0}$ (depending on norm/architecture control of $\mathcal{F}_{K}$ but not on $n$) such that
@@
Apply Lemma B.15 to $\mathcal{H}_{K,S}$ to obtain (18). For the complexity term, write $\mathcal{H}_{K,S}=\ell\circ\mathcal{G}_{K,S}$. By Lemma B.16, $\mathfrak{R}_{M}(\mathcal{H}_{K,S})\leq L_{\ell}\mathfrak{R}_{M}(\mathcal{G}_{K,S})$, and by B.17, $\mathfrak{R}_{M}(\mathcal{G}_{K,S})\leq C_{0}(L_{f}R/\sqrt{M})\sqrt{K}$. Combining yields (19) with $C=C_{0}$ (or another universal constant).
@@ Appendix F Full Metrics for Main Comparison
Table 3: Additional metrics for main comparison (miRAW half-split).
@@
TargetNet (reproduced) 0.7226 0.6560 0.9361 0.5091
Max pooling (ours) 0.8049 0.7822 0.8453 0.7646
PAIR-Former (ours) 0.9744 0.9886 0.9598 0.9890
@@ Appendix H Additional runtime profiling details
This appendix documents the profiling protocol behind Figures 3 and 4. We measure online inference cost for three pipelines: BR-MIL_online (cheap scan + STSelector + expensive encode on $K$ + Set Transformer), TargetNet_like_online (window/CTS scoring + pooling; budget-independent), and a heavier Naive_online variant that performs more per-candidate computation before aggregation.

### v2

@@ Conclusion.
Assume B.16, B.17, B.18, B.19 and B.24. Fix any selector $S$ (deterministic or conditioned upon), and consider $\mathcal{H}_{K,S}$. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the draw of $\mathcal{D}$, for all $f\in\mathcal{F}_{K}$,
@@
If additionally B.20 holds with $|\mathcal{S}_{K}|=N_{S}<\infty$, then with probability at least $1-\delta$, uniformly for all $(f,S)\in\mathcal{F}_{K}\times\mathcal{S}_{K}$,
@@
Assume B.18. Let $\mathcal{G}$ be a class of real-valued functions on bags, and define $\ell\circ\mathcal{G}:=\{(B,Y)\mapsto\ell(Y,g(B)):\ g\in\mathcal{G}\}$. Then
@@
Assumption B.24 (Capacity control of $\mathcal{F}_{K}$ – derived result).
@@
By Theorem B.13 (derived in Appendix B.3), there exists a constant $C_{0}$ (depending on norm/architecture control of $\mathcal{F}_{K}$ but not on $n$) such that
@@
Apply Lemma B.22 to $\mathcal{H}_{K,S}$ to obtain (20). For the complexity term, write $\mathcal{H}_{K,S}=\ell\circ\mathcal{G}_{K,S}$. By Lemma B.23, $\mathfrak{R}_{M}(\mathcal{H}_{K,S})\leq L_{\ell}\mathfrak{R}_{M}(\mathcal{G}_{K,S})$, and by B.24 (derived in Appendix B.3), $\mathfrak{R}_{M}(\mathcal{G}_{K,S})\leq C_{0}(L_{f}R/\sqrt{M})\sqrt{K}$. Combining yields (21) with $C=C_{0}$ (or another universal constant).
@@
B.5 Practical Guidance for Budget Selection
We now discuss how the theoretical bounds can inform the practical selection of the budget $K$.
Bias-variance tradeoff in $K$.
Theorems B.8 and B.21 reveal a fundamental bias-variance tradeoff controlled by $K$:
  • Approximation error (bias): By Theorem B.8, the gap between the full-information predictor and the budgeted predictor is bounded by $2L_{f}R\cdot\varepsilon_{K}$, which decreases as $K$ increases (more instances are covered, reducing truncation error).
  • Generalization error (variance): By Theorem B.21, the dominant generalization term scales as $C_{\text{arch}}\cdot R\sqrt{K}/\sqrt{M}$, which increases with $K$ (larger input space leads to higher model capacity and potential overfitting).
Optimal $K$ in theory.
The total excess risk (approximation + generalization) is bounded by
$\begin{split}\text{ExcessRisk}(K)&\leq\underbrace{2L_{f}R\cdot\varepsilon_{K}}_{\text{approximation (decreasing in }K\text{)}}\\ &+\underbrace{C_{\text{arch}}\cdot\frac{R\sqrt{K}}{\sqrt{M}}}_{\text{generalization (increasing in }K\text{)}}\\ &+\text{lower-order terms}.\end{split}$ (23)
If the uncovered influence mass decays as $\varepsilon_{K}\leq CK^{1-\alpha}$ for some $\alpha>1$ (Assumption B.6), then the optimal $K^{*}$ balances the two terms:
$K^{1-\alpha}\sim\frac{\sqrt{K}}{\sqrt{M}}\quad\Longrightarrow\quad K^{*}\sim M^{\frac{1}{2\alpha-1}}.$ (24)
For example:
  • If $\alpha=2$ (fast tail decay): $K^{*}\sim M^{1/3}$. With $M\approx 5000$ training pairs, this gives $K^{*}\approx 17$.
  • If $\alpha=3/2$ (moderate tail decay): $K^{*}\sim M^{1/2}$. With $M\approx 5000$, this gives $K^{*}\approx 71$.
  • If $\alpha=3$ (very fast tail decay): $K^{*}\sim M^{1/5}$. With $M\approx 5000$, this gives $K^{*}\approx 5.5$.
Remark B.25 (Consistency with empirical $K^{*}$).
Our empirical operating point $K^{*}=64$ is consistent with the moderate tail decay regime ($\alpha\approx 3/2$), which aligns with the heavy-tailed CTS pool structure observed in miRNA targeting (many low-influence candidates, a few high-influence ones; see Figure 2). The performance saturation observed in Fig. 3 around $K\geq 64$ is consistent with the approximation error becoming negligible at this budget.
Practical recommendation.
While the exact optimal $K^{*}$ depends on the unknown tail parameter $\alpha$, the theory suggests the following practical guidelines:
  1. 1.
    Start with $K\approx\sqrt{M}$ as a reasonable default (this corresponds to the moderate tail decay regime).
  2. 2.
    Sweep $K$ on validation data to find the empirical optimum, using the theory-predicted monotone improvement in approximation and potential overfitting at large $K$ as guidance.
  3. 3.
    Monitor the approximation-generalization tradeoff: if increasing $K$ improves training loss but not validation loss, the generalization term is dominating and $K$ should be reduced.
B.6 Tightness of the $\sqrt{K}$ Bound
We briefly discuss the tightness of the $\sqrt{K}$ dependence in the Rademacher complexity bound.
Proposition B.26 (Lower bound on Rademacher complexity).
For any permutation-invariant function class $\mathcal{F}_{K}$ that includes linear functions of the form $f(X)=\langle w,\bar{x}\rangle$ where $\bar{x}=\frac{1}{K}\sum_{i=1}^{K}z_{i}$ is the mean token, the Rademacher complexity satisfies
$\mathfrak{R}_{M}(\mathcal{F}_{K})\geq c\cdot\frac{R}{\sqrt{M}}\cdot\frac{1}{\sqrt{K}},$ (25)
for a constant $c>0$ depending on the token dimension $d_{z}$.
Proof sketch.
The mean pooling function $f(X)=\langle w,\bar{x}\rangle$ with $\|w\|_{2}\leq 1$ has Rademacher complexity $\Omega(R/(\sqrt{K}\cdot\sqrt{M}))$ because the mean $\bar{x}$ has norm $\|\bar{x}\|_{2}\leq R/\sqrt{K}$ (by concentration of the mean of bounded vectors). ∎
Remark B.27 (Gap between upper and lower bounds).
The upper bound is $O(R\sqrt{K}/\sqrt{M})$ (Theorem B.13) and the lower bound is $\Omega(R/(\sqrt{K}\sqrt{M}))$ (Proposition B.26). The gap of $K$ between upper and lower bounds reflects the difference between the full Set Transformer (which can model complex interactions among $K$ tokens) and simple mean pooling (which only uses the average). For practical Set Transformers, the true complexity likely lies between these extremes, and the upper bound $O(\sqrt{K})$ is the relevant quantity for generalization control.
@@ Appendix E Loss Functions and Distillation Details
Ablation of $\mathcal{L}_{\mathrm{rel}}$.
To assess the contribution of relational distillation, we ablate it ($\beta_{\mathrm{rel}}=0$) and retrain both Stage 2 and Stage 3. Downstream Stage 3 performance is virtually unchanged (F1 $0.975{\pm}0.002$ vs. $0.975{\pm}0.002$; PR-AUC $0.989{\pm}0.001$ vs. $0.989{\pm}0.001$), indicating that cheap encoder quality at the current dataset scale is dominated by the supervised and logit distillation terms. We retain $\mathcal{L}_{\mathrm{rel}}$ for completeness of the distillation formulation, but do not claim a measurable benefit at the current dataset scale.
@@ Appendix F Full Metrics for Main Comparison
[…]
@@
[…]
@@ Appendix H Additional runtime profiling details
This appendix documents the profiling protocol behind Figures 4 and 5. We measure online inference cost for three pipelines: BR-MIL_online (cheap scan + STSelector + expensive encode on $K$ + Set Transformer), TargetNet_like_online (window/CTS scoring + pooling; budget-independent), and a heavier Naive_online variant that performs more per-candidate computation before aggregation.
@@
Appendix I Split Sensitivity Analysis
To verify that our results are not an artifact of the specific train/test partition, we evaluate PAIR-Former under two additional splits of the 10 released miRAWtest subsets, each with R=3R{=}3 seeds $\{2020,2025,2026\}$. All hyperparameters, architecture, and training protocol are held constant.
[…]
Split Train / Test F1@0.5 PR-AUC
[…]
Performance is highly consistent: PR-AUC varies by only ${\sim}0.006$ and F1 by ${\sim}0.004$ across three fundamentally different partitioning strategies (the original non-overlapping split, a consecutive block boundary, and a maximally interleaved assignment), confirming that the BR-MIL pipeline is not sensitive to the particular train/test partition.
Appendix J Data Overlap Statistics
We provide detailed overlap statistics to address disjointness concerns raised by reviewers.
Table 6: Overlap between Stage 1–2 CTS training and Stage-3 test data
Level CTS Data Stage-3 Test Overlap % of Test
mRNA transcripts 6,081 2,534 1,745 68.9%
miRNA IDs 429 667 361 54.1%
Exact (miRNA, mRNA) pairs 24,745 3,270 76 2.3%
Table 7: Overlap within Stage-3 train and test splits
Category Train Test Overlap % of Test
Positive pairs 2,718 2,722 60 2.2%
Negative pairs 548 548 548 100%
Total unique 3,266 3,270 608 18.6%
Note on overlap interpretation.
The 60 unique overlapping positive pairs correspond to 61 test examples (one pair appears twice in the test set). The 100% negative overlap stems from the construction of miRAWtest: only 548 unique negative pairs exist across all 10 subsets, which are reused. This affects all methods evaluated under this benchmark, not only ours.